The Lunder Company has total assets of $22,542,000, current liabilities of $2,547,000, and long-term liabilities of $7,410,000. The firm has 1,000,000 shares of common stock outstanding. Compute the firm’sbook value of equity per share.

Referring back to the previous question, if Lunder’s net income equals $1,600,000 and Lunder has a P/E ratio of 15.0 (note that P/E ratio = stock price per share divided by earnings per share), what is the current price of the stock?
Answer #1

Solution:

Shareholder equity = Total assets - Current liabilities - long term liabilities

Shareholder equity = $2,25,42,000 - $25,47,000 - $74,10,000= $1,25,85,000

Number of shares = 10,00,000

Book value per share = 1,25,85,000/1000000= $12.585 rounded to $12.59

Net income = $16,00,000

PE ratio = 15

Earning per share = Net income / number of shares

Earning per share = 16,00,000/1000000= =$1.60

Current price = PE ratio * Earning per share

Current price = 15*1.60= $24 per share